# Unified Domain Adaptive Semantic Segmentation

### v2

@@ IV-B1 Comparison With State-of-the-Arts of UDA-VSS
| DAFormer[24] | Trans. | 95.7 | 70.2 | 89.4 | 53.5 | 48.1 | 49.6 | 55.8 | 59.4 | 89.9 | 47.9 | 92.5 | 72.2 | 44.7 | 92.3 | 74.5 | 78.2 | 65.1 | 55.9 | 61.8 | 68.3 |
| SePiCo[23] | Trans. | 96.7 | 76.7 | 89.7 | 55.5 | 49.5 | 53.2 | 60.0 | 64.5 | 90.2 | 50.3 | 90.8 | 74.5 | 44.2 | 93.3 | 77.0 | 79.5 | 63.6 | 61.0 | 65.3 | 70.3 |
| FREDOM[77] | Trans. | 96.7 | 74.8 | 90.9 | 58.1 | 49.0 | 57.5 | 63.4 | 71.4 | 91.6 | 52.1 | 94.4 | 78.4 | 53.1 | 94.1 | 83.9 | 85.2 | 72.5 | 62.8 | 68.9 | 73.6 |
| QuadMix | Trans. | 97.5 | 80.9 | 91.6 | 62.3 | 57.6 | 58.2 | 64.5 | 71.2 | 91.7 | 52.3 | 94.3 | 80.0 | 55.9 | 94.6 | 86.3 | 90.5 | 82.3 | 65.1 | 68.1 | 76.1 |
@@
| QuadMix | DL-V2 | 88.5 | 52.9 | 87.1 | 3.7 | 1.4 | 56.2 | 62.7 | 59.2 | 87.2 | 89.0 | 79.1 | 55.8 | 87.9 | 66.7 | 58.1 | 71.2 | 62.6 | 72.3 |
| DAFormer[24] | Trans. | 84.5 | 40.7 | 88.4 | 41.5 | 6.5 | 50.0 | 55.0 | 54.6 | 86.0 | 89.8 | 73.2 | 48.2 | 87.2 | 53.2 | 53.9 | 61.7 | 60.9 | 67.4 |
| SePiCo[23] | Trans. | 87.0 | 52.6 | 88.5 | 40.6 | 10.6 | 49.8 | 57.0 | 55.4 | 56.8 | 86.2 | 75.4 | 52.7 | 92.4 | 78.9 | 53.0 | 62.6 | 64.3 | 71.4 |
| FREDOM[77] | Trans. | 89.4 | 50.8 | 89.3 | 48.8 | 9.3 | 57.3 | 65.1 | 60.1 | 89.9 | 93.7 | 79.4 | 51.6 | 90.5 | 66.0 | 62.3 | 68.1 | 67.0 | 73.6 |
| QuadMix | Trans. | 88.1 | 51.2 | 88.9 | 46.7 | 7.9 | 58.6 | 64.7 | 63.7 | 88.1 | 93.9 | 81.3 | 56.6 | 90.3 | 66.9 | 66.8 | 66.0 | 67.5 | 74.3 |
@@
IV-B2 Comparison With State-of-the-Arts of UDA-ISS
For UDA-ISS, we replace spatio-temporal QuadMix with spatial QuadMix, and substitute optical guided spatio-temporal aggregation with spatial aggregation for fine-grained alignment. Technically, we shift elements of and as defined in Section III-B, and remove the reliance on temporal clues that are not present in image scenarios. Additionally, target pseudo-labels are generated using a momentum network [72].
Table III and Table IV compare our method with existing UDA-ISS approaches. Impressively, we maintain superiority on two challenging benchmarks: GTAV Cityscapes and Synthia Cityscapes. We compare with recent advanced CNN-based methods [20, 17, 19, 5, 73, 21, 76, 22, 23, 77] and transformer-based methods [24, 23, 77]. Notably, DACS [19] proposes one-way mixing, which is followed by subsequent works [24, 23, 77] combing contrastive learning or long-tail learning. Other works [21, 22] employ two-way mixing, with the latter integrating style transfer. However, as analyzed above, they overlooked distinct point attributes within each domain, leading to fragmentary and insufficient domain gap bridging. Additionally, their focus on pixel-level mixing alone can result in feature inconsistencies, hindering effective knowledge transfer. In contrast, we address these challenges comprehensively and obtain remarkable results.
GTAV Cityscapes. As shown in Table III, our method achieves mIOUs of 66.8 with DeepLab-V2 and 76.1 with transformer networks. Notably, QuadMix (Trans.) more effectively extracts spatial features from long-tail categories. Specifically, for the train class, our method achieves an IOU of 82.3, surpassing that of FREDOM [77] (72.5), which is tailored for long-tail learning. Notably, QuadMix (DL-V2) underperforms for the train class due to the limited feature extraction capability of CNNs. We believe that more frequent resampling of such categories could improve the results. Generally, our method boosts the performance of most categories, exhibiting robust improvement.
Synthia Cityscapes. In Table IV, we present results for 13-category (mIOU*) and 16-category (mIOU) adaptation, following [23, 22], etc. BDM [21] only studies 13-category adaption. Our method consistently outperforms previous works and is well-adapted for transformers and particularly CNN networks on this benchmark, validating its effectiveness.
@@ IV-C Ablation Study
In this section, we conduct ablation studies to validate quad-directional mixing and flow-guided feature aggregation, using the Deeplabv2 backbone for UDA-VSS tasks by default.
| ID | SSL | V-tmpl | F-tmpl | Agg. | Warm | mIOU (Gain) |
| ✓ | 52.01 | |||||
| 1 | ✓ | ✓ | 55.71 (+3.70) | |||
| 2 | ✓ | ✓ | ✓ | 57.25 (+5.24) | ||
| 3 | ✓ | ✓ | 54.01 (+2.00) | |||
| 4 | ✓ | ✓ | ✓ | ✓ | 59.62 (+7.61) | |
| 5 | ✓ | ✓ | ✓ | ✓ | 57.95 (+5.95) | |
| 6 | ✓ | ✓ | ✓ | 55.97 (+3.96) | ||
| 7 | ✓ | ✓ | ✓ | ✓ | 60.26 (+8.25) | |
| 8 | ✓ | ✓ | ✓ | ✓ | ✓ | 61.48 (+9.47) |
@@ IV-C1 Ablation Study on Each Component
In Table V, we conduct ablation experiments to verify the effectiveness of three key components: video patch templates (V-tmpl) mixing, feature-level templates (F-tmpl) mixing, and optical-guided video feature aggregation (Agg.). Additionally, we also study the impact of the warm-up mentioned in Table I and Table II. For a fair comparison, we take the results of TPS [37] without random scale augmentation as the baseline (Exp. ID ).
It can be concluded that: (i) Video patch templates (V-tmpl) result in a 3.70 mIOU gain in Exp. ID 1; (ii) As for feature-level template mixing (F-tmpl) across two enhanced inter-mixed domains, Exp. ID 2 further incorporating F-tmpl demonstrates a 1.54 mIOU gain. In addition, Exp. 8 outperforms Exp. 7 by 1.22, also validating the efficacy of “F-tmpl”; (iii) In another line, Exp. 3 introduces the optical-guided video feature aggregation (Agg.), resulting in growth to 54.01; (iv) Furthermore, Exp. ID 4 combines “V-tmpl”, “F-tmpl”, and “Agg.”, achieving an mIOU of 59.62, surpassing the state-of-the-art TPL-SFC [38] mIOU of 55.3 by 4.32; (v) Moreover, we also employ the warm-up setup in Exp. ID 5 to Exp. ID 8 for model parameter initialization, exceeding the state-of-the-art CMOM of 56.31 with a warm-up stage by 5.17, resulting in the highest mIOU of 61.48, a remarkable 9.47 increase compared to the baseline Exp. ID .
Further comparison. To valid the video patch templates (V-tmpl) in Exp. ID 1, we perform a variant using image patch templates (I-tmpl) akin to [21, 19]. Thus, the model input is also substituted with and in Eq. (22). It is observed that with “I-tmpl”, the mIOU dropped by 1.37 compared with Exp. ID 1, which suggests that “V-tmpl” enhances the video adaptation more effectively.
Qualitative results of each component. In Fig. 8, we present segmentation results of three adjacent frames on Synthia-Seq Cityscapes-Seq, following experiment settings corresponding to Exp. ID 0*, ID 2, ID 3, ID 5, ID 8 in Table V. These results demonstrate the incremental integration of each component. The segmentation results progressively improve from left to right, particularly for categories like sign and car. Notably, the challenging issue with a person on the far left within the frame is effectively addressed by integrating the proposed key modules (Exp. ID 8).
@@ IV-C2 Ablation Study on Quad-Directional Mixing
It can be observed: (i) In Exp. ID 1, our setting outperformed the CMOM [39] by 1.80 in mIoU, which also follows a source to target pasting paradigm. We attribute this improvement to the proposed flow-guided feature aggregation for domain alignment; (ii) The construction of intra/inter-mixed video domains always yields better results than solely one-way mixed domains. It is because the model can learn consistent spatio-temporal features with diverse domain contexts, with the feature-level mixing across two inter-mixed domains alleviating feature inconsistencies; (iii) Utilizing source domain templates commonly outperforms using target domain templates with higher predictive confidence. The overall mIOU exhibits further improvement when incorporating templates from both domains; (iv) More intriguingly, compared with incrementally combining various directional mixing as in Exp. ID 11, we not only reduce computational load significantly as shown in Table XI, but also achieve a mIoU gain of 0.92 to 61.48 by fully utilizing the enhanced quad-mixed domains for comprehensive domain bridging (Exp. ID 12).
It’s noteworthy that regardless of the intra/inter-mixing strategy adopted, the results of ablation studies in Table VI all surpass previous state-of-the-art methods CMOM [39] and TPL [38], which highlights the effectiveness of our QuadMix, a novel and unified perspective that has been largely overlooked by prior works [39, 37, 19, 73, 21], etc.
Feature visualization for quad-directional mixing paths. In Fig. 7, following experimental setings in Table VI, we present t-SNE [43] feature visualization for pure two domains (Exp. ID 0*), our intra-mixed domains (Exp. ID 6), one-way mixing (Exp. ID 1), two-way mixing (Exp. ID 3), and the proposed QuadMix (Exp. ID 12) from left to right. Notably, the number of points in Fig. 7(b)(d)(e) are equal. It can be found that quad-directional mixing effectively generalizes intra-domain features and enhances inter-domain features, addressing distinct point attributes within each domain. By sufficiently bridging domain gaps, it promotes dense knowledge transfer from the source domain to the target domain.
@@
The effect of category selection on video patch template generation. The video patch templates used for quad-mixing involve patch regions corresponding to the randomly selected two categories from the source or target domain. We investigated the impact of category selection on generating video patch templates. Commonly, drawing on the concept from panoramic segmentation [82], objects are often categorized into “things” (e.g., sign, car) and “stuff” (e.g., sky, road). Therefore, in Table VII, we examined the results when the category selection comprised only “things” or “stuff”, as well as movable (e.g., person, car) or stationary (e.g., sign, light) objects. The experimental findings indicate that “things” and movable categories yield better results than “stuff” and stationary categories, and the model performs best when the category selection encompasses all categories. It highlights the versatility of our method across different categories.
@@
IV-C3 Ablation Study on Flow-Guided Feature Aggregation for Domain Alignment
In this paper, to further address spatio-temporal gaps within videos, we propose the optical guided video feature aggregation module for fine-grained alignment. When we exclusively substitute our fine-grained aggregation with treating video features as a whole [73], the mIOUs on two benchmarks decreased by 3.24 and 2.54, respectively, as shown in Table VIII. It demonstrates the effectiveness of fine-grained aggregation in compressing video features into a more compact distribution, thereby reducing domain discrepancy with fine-grained alignment. In addition, we also investigate the advantage of image enhancement in Eq. (22) and Eq. (24) for the target domain. Note that our contribution focuses on enhancing temporal coherence for domain alignment, which alignment loss is optimal to minimize discrepancies in aggregated features across domains is another research branch and goes beyond the main consideration of this work.
IV-D Further Analysis
In this section, we conduct further analysis using UDA-VSS benchmarks with the Deeplabv2 backbone by default.

### v3

@@ IV-A3 Implementation Details
| DAFormer[1] | ViT | 95.7 | 70.2 | 89.4 | 53.5 | 48.1 | 49.6 | 55.8 | 59.4 | 89.9 | 47.9 | 92.5 | 72.2 | 44.7 | 92.3 | 74.5 | 78.2 | 65.1 | 55.9 | 61.8 | 68.3 |
| SePiCo[23] | ViT | 96.7 | 76.7 | 89.7 | 55.5 | 49.5 | 53.2 | 60.0 | 64.5 | 90.2 | 50.3 | 90.8 | 74.5 | 44.2 | 93.3 | 77.0 | 79.5 | 63.6 | 61.0 | 65.3 | 70.3 |
| FREDOM[67] | ViT | 96.7 | 74.8 | 90.9 | 58.1 | 49.0 | 57.5 | 63.4 | 71.4 | 91.6 | 52.1 | 94.4 | 78.4 | 53.1 | 94.1 | 83.9 | 85.2 | 72.5 | 62.8 | 68.9 | 73.6 |
| QuadMix | ViT | 97.5 | 80.9 | 91.6 | 62.3 | 57.6 | 58.2 | 64.5 | 71.2 | 91.7 | 52.3 | 94.3 | 80.0 | 55.9 | 94.6 | 86.3 | 90.5 | 82.3 | 65.1 | 68.1 | 76.1 |
@@
| QuadMix | DL-V2 | 88.5 | 52.9 | 87.1 | 3.7 | 1.4 | 56.2 | 62.7 | 59.2 | 87.2 | 89.0 | 79.1 | 55.8 | 87.9 | 61.7 | 58.1 | 71.2 | 62.6 | 72.3 |
| DAFormer[1] | ViT | 84.5 | 40.7 | 88.4 | 41.5 | 6.5 | 50.0 | 55.0 | 54.6 | 86.0 | 89.8 | 73.2 | 48.2 | 87.2 | 53.2 | 53.9 | 61.7 | 60.9 | 67.4 |
| SePiCo[23] | ViT | 87.0 | 52.6 | 88.5 | 40.6 | 10.6 | 49.8 | 57.0 | 55.4 | 56.8 | 86.2 | 75.4 | 52.7 | 92.4 | 78.9 | 53.0 | 62.6 | 64.3 | 71.4 |
| FREDOM[67] | ViT | 89.4 | 50.8 | 89.3 | 48.8 | 9.3 | 57.3 | 65.1 | 60.1 | 89.9 | 93.7 | 79.4 | 51.6 | 90.5 | 66.0 | 62.3 | 68.1 | 67.0 | 73.6 |
| QuadMix | ViT | 88.1 | 51.2 | 88.9 | 46.7 | 7.9 | 58.6 | 64.7 | 63.7 | 88.1 | 93.9 | 81.3 | 56.6 | 90.3 | 66.9 | 66.8 | 66.0 | 67.5 | 74.3 |
IV-B Experimental Results
IV-B1 Comparison With State-of-the-Arts of Video UDA-SS
In Table I and Table II, we compare our method with the state-of-the-art methods on Synthia-Seq Cityscapes-Seq and VIPER Cityscapes-Seq benchmark. To the best of our knowledge, recent methods [16, 44, 17, 18, 19, 20, 2] are the most superior works in video UDA-SS, all employing CNN-based architectures. Notably, for the first time, we explore transformer-based architectures for video UDA-SS. Moreover, we also compare several variants to video UDA-SS that are initially designed for image UDA-SS, including works [53, 54, 55, 13, 56, 57, 58, 59, 10, 21, 60]. Technically, following the previous video UDA-SS works [16, 44, 17, 2], we replace their image segmentation backbones with a video segmentation network (denoted by †). Additionally, “Adv.” signifies methods that employ adversarial learning, and “SSL” denotes self-supervised learning. “Warm” denotes initializing parameters of the DeepLab-V2 model with a warm-up model, whereas others use ResNet-101 pre-trained on ImageNet. “ViT” refers to the transformer backbone.
Synthia-Seq Cityscapes-Seq. In Table I, we present experimental results comparing our method with state-of-the-art techniques on the Synthia-Seq Cityscapes-Seq. For clarity, the best per-category IOU and mIOU results are in bold, while suboptimal results are underlined. We observed that: (i) Our method is highly compatible with both CNN and transformer-based architectures. The proposed CNN-based Quadmix significantly outperforms the previous best CMOM [2] by 5.2 mIOU. When equipped with a transformer, we further boost mIOU by 5.7, achieving the highest result of 67.2; (ii) For CNN-based QuadMix, without a warm-up model, we achieved 59.6 mIOU, which is 4.3 higher than TPL-SFC [19] by a large margin. When using a warm-up model, we outperform the CMOM, which relies heavily on the warm-up model in [16] for both model initialization and pseudo-label generation; (iii) Adversarial learning methods can facilitate domain adaptation but are less improving compared to self-supervised methods; (iv) Our method boosts the IOU for most categories. It can be attributed to our comprehensive quad-directional mixing, which bridges the feature gap effectively and addresses the intra-domain discontinuity and feature inconsistencies, as well as flow-guided video feature aggregation for cross-domain alignment.
Viper Cityscapes-Seq. The domain gap between Viper and Cityscapes-Seq is much larger than that of Synthia-Seq and Cityscapes-Seq. As shown in Table II, our proposed QuadMix also achieves state-of-the-art results. We can conclude that: (i) For CNN-based QuadMix, we achieve 56.2 mIOU with a warm-up model and 54.9 mIoU without it; furthermore, we obtain 58.2 mIOU when we employ transformer-based network. All our results significantly outperform previous methods; (ii) Recent work I2VDA [44] focuses exclusively on transferring source spatial information to target domain videos, resulting in a target domain mIOU of 51.2. This is inferior to our model, which comprehensively investigates both spatial and temporal level domain shifts. (iii) Variants of image domain adaptation methods generally underperform on target videos compared to video UDA-SS methods, likely due to inadequate utilization of crucial temporal information.
@@ IV-B1 Comparison With State-of-the-Arts of Video UDA-SS
IV-B2 Comparison With State-of-the-Arts of Image UDA-SS
For image UDA-SS, we replace spatio-temporal QuadMix with spatial QuadMix, and substitute flow-guided spatio-temporal feature aggregation with spatial feature aggregation. Technically, we shift elements of and for images defined in Section III-B, and remove the reliance on temporal cues that are not present in image scenarios. Additionally, target pseudo-labels are generated using a momentum network [51].
Table III and Table IV compare our method with existing image UDA-SS approaches. Impressively, we maintain superiority on two challenging benchmarks: GTAV Cityscapes and SYNTHIA Cityscapes. We compare with recent advanced CNN-based methods [60, 58, 21, 10, 3, 5, 66, 22, 23, 67] and transformer-based methods [1, 23, 67]. Notably, DACS [21] proposes one-way mixing, which is followed by subsequent works [1, 23, 67] combing contrastive learning or long-tail learning. Other works [5, 22] employ two-way mixing, with the latter integrating style transfer. However, as analyzed above, they overlooked the intra-domain discontinuity within each domain, leading to fragmented and insufficient domain gap bridging. Additionally, their focus on pixel-level mixing alone can result in feature inconsistencies, hindering effective knowledge transfer. In contrast, we address these challenges comprehensively and obtain remarkable results.
GTAV Cityscapes. As shown in Table III, our method achieves mIOUs of 66.8 with DeepLab-V2 and 76.1 with transformer networks. Notably, QuadMix (ViT) more effectively extracts spatial features from long-tail categories. Specifically, for the train class, our method achieves an IOU of 82.3, surpassing that of FREDOM [67] (72.5), which is tailored for long-tail learning. Notably, QuadMix (DL-V2) underperforms for the train class due to the limited feature extraction capability of CNNs. We believe that more frequent resamples for models to learn such categories could improve the results. Generally, our method boosts the performance of most categories, exhibiting robust improvement.
SYNTHIA Cityscapes. In Table IV, we present results for 13-category (mIOU*) and 16-category (mIOU) adaptation, following [23, 22], etc. BDM [5] only studies 13-category adaption. Note that QuadMix (DL-V2) has already outperformed DAFormer (ViT) by 1.7 in 13-category adaptation and 4.9 in 16-category adaptation, with further improvements of 4.9 and 2.0, respectively, when equipped with ViT. Overall, our method consistently demonstrates superior performance and is well-adapted for transformers and CNNs.
For a qualitative comparison with previous works, please refer to the supplementary material.
@@ IV-C Ablation Study
Given that an image is essentially a special case of a video without the temporal dimension, we opted for the more general video UDA-SS for ablation studies, using the DeepLab-V2 backbone by default.
| ID | V-template | F-template | Agg. | Warm | mIOU (Gain) |
| 52.01 | |||||
| 1 | ✓ | 55.71 (+3.70) | |||
| 2 | ✓ | ✓ | 57.25 (+5.24) | ||
| 3 | ✓ | 54.01 (+2.00) | |||
| 4 | ✓ | ✓ | ✓ | 59.62 (+7.61) | |
| 5 | ✓ | ✓ | ✓ | 57.95 (+5.94) | |
| 6 | ✓ | ✓ | 55.97 (+3.96) | ||
| 7 | ✓ | ✓ | ✓ | 60.26 (+8.25) | |
| 8 | ✓ | ✓ | ✓ | ✓ | 61.48 (+9.47) |
@@ IV-C1 Ablation Study on Each Component
In Table V, we conduct ablation studies to verify the effectiveness of three key components: video patch template (V-template) mixing, feature-level template (F-template) mixing, and optical-guided video feature aggregation (Agg.). Additionally, we also study the impact of the warm-up mentioned in Table I and Table II. For a fair comparison, we take the result of TPS [17] without random scale augmentation as the baseline (Exp. ID ).
It can be concluded that: (i) Video patch template (V-template) result in a 3.70 mIOU gain in Exp. ID 1; (ii) As for feature-level template mixing (F-template) across two enhanced quad-mixed domains, Exp. ID 2 further incorporates the F-template and demonstrates a 1.54 mIOU gain. In addition, Exp. 8 outperforms Exp. 7 by 1.22, also validating the efficacy of “F-template”; (iii) In another line, Exp. 3 introduces the optical-guided video feature aggregation (Agg.), resulting in growth to 54.01; (iv) Furthermore, Exp. ID 4 combines “V-template”, “F-template”, and “Agg.”, achieving an mIOU of 59.62, surpassing the state-of-the-art TPL-SFC [19] mIOU of 55.3 by 4.32; (v) Moreover, we also employ the warm-up setup in Exp. ID 5 to Exp. ID 8 for model parameter initialization, exceeding the state-of-the-art CMOM of 56.31 with a warm-up stage by 5.17, resulting in the highest mIOU of 61.48, a remarkable 9.47 increase compared to the baseline Exp. ID .
Further comparison. To valid the video patch templates (V-template) in Exp. ID 1, we perform a variant using image patch templates (I-template) akin to [5, 21]. Thus, the model input is also substituted with and in Eq. (22). It is observed that with “I-template”, the mIOU dropped by 1.37 compared with Exp. ID 1, which suggests that “V-template” enhances the video adaptation more effectively.
Qualitative results of each component. In Fig. 8, we present segmentation results of three adjacent frames on Synthia-Seq Cityscapes-Seq, following experiment settings corresponding to Exp. ID 0*, ID 2, ID 3, ID 5, ID 8 in Table V. These results demonstrate the incremental integration of each component. The segmentation results progressively improve from left to right, particularly for categories like sign and car. Notably, the challenging issue with a person on the far left within the frame is effectively addressed by integrating the proposed key components (Exp. ID 8).
@@ IV-C2 Ablation Study on Quad-Directional Mixing
It can be observed: (i) In Exp. ID 1, our setting outperformed the CMOM [2] by 1.80 in mIoU, which also follows a source to target pasting paradigm. We attribute this improvement to the proposed flow-guided feature aggregation for domain alignment; (ii) The construction of intra/inter-mixed video domains always yields better results than solely one-way mixed domains. It is because the model can learn consistent spatio-temporal features with diverse domain contexts, with the feature-level mixing across quad-mixed domains alleviating feature inconsistencies; (iii) Utilizing source domain templates commonly outperforms using target domain templates with higher predictive confidence. The overall mIOU exhibits further improvement when incorporating templates from both domains; (iv) More intriguingly, compared with incrementally combining various directional mixing as in Exp. ID 11, we not only reduce computational load significantly as shown in Table XI, but also achieve a mIoU gain of 0.92 to 61.48 by fully utilizing the enhanced quad-mixed domains for comprehensive domain bridging (Exp. ID 12).
It’s noteworthy that regardless of the intra/inter-mixing strategy adopted, the results of ablation studies in Table VI all surpass previous state-of-the-art methods CMOM [2] and TPL [19], which highlights the effectiveness of our novel and unified perspective that has been largely overlooked by prior works [2, 17, 21, 3, 5], etc.
Feature visualization for quad-directional mixing paths. In Fig. 9, following experimental setings in Table VI, we present t-SNE [50] feature visualization for pure two domains (Exp. ID 0*), our intra-mixed domains (Exp. ID 6), one-way mixing (Exp. ID 1), two-way mixing (Exp. ID 3), and the proposed QuadMix (Exp. ID 12) from left to right. Notably, the number of points in Fig. 9(b)(d)(e) are equal. It can be found that quad-directional mixing effectively generalizes intra-domain features and enhances inter-domain features, addressing the insufficient feature gap bridging across distinct domains. By bridging domain gaps sufficiently, it promotes dense knowledge transfer from the source to the target domain.
@@
The effect of category selection on video patch template generation. The video patch templates used for quad-mixing involve patch regions corresponding to the randomly selected two categories from the source or target domain. We investigated the impact of category selection on generating video patch templates. Drawing on the concept from panoramic segmentation [72], objects are often categorized into “things” (e.g., sign, car) and “stuff” (e.g., sky, road). Therefore, in Table VII, we examined the results when the category selection comprised only “things” or “stuff”. Moreover, we also explore the movable (e.g., person, car) or stationary (e.g., sign, light) objects. The experimental findings indicate that “things” and movable categories yield better results than “stuff” and stationary categories, and the model performs best when the category selection encompasses all categories, which highlights the versatility of our method across different categories.
@@
IV-C3 Ablation Study on Flow-Guided Spatio-Temporal Feature Aggregation Module
In this paper, to further address spatio-temporal gaps across video domains, we propose the optical flow-guided video feature aggregation module for fine-grained alignment. When exclusively substituting our fine-grained aggregation with treating video features as a whole [3], the mIOUs on two benchmarks decreased by 3.24 and 2.54, respectively, as shown in Table VIII. It demonstrates the effectiveness of our method in compressing video features across domains into a more aligned distribution, thereby reducing the domain discrepancy with precise alignment. In addition, we also investigate the effect of image enhancement in Eq. (22) and Eq. (24) for target domain samples.
Note that our contribution focuses on enhancing spatio-temporal coherence of aggregated features for domain alignment, which alignment loss is optimal to minimize discrepancies of aggregated features across domains is another research branch and goes beyond the main consideration of this work.